the cost for the manufacturing of a cd varies inversely with the number of cd's made. if the cost is $2.50 per cd when 10,000 are made, then what is the cost per cd when 100,000 are made?
Answer by Theo(13342)

the formula for inverse variation is:

y = k/x

k is the constant of variation.

let y = cost per cd
let x = number of cd

y = k/x becomes:

2.50 = k / 10,000

Solve for k to get k = 10,000 * 2.50 = 25,000

Since k is a constant, it will not change.

cost when 100,000 cd are made would therefore be:

y = 25,000 / 100,000 = .25 per cd.